about ages of people in apartodar regon A have te Rotowing ages , in years. Complete parts (a) tough (d)b a. Determine the range of these ages. The range is- Simplity your answer.) the sample standard deviation of these ages by using the defining formiula s b. Find sample size and x is the sample mean. The standard deviation is (Round to one decimal place as needed.) where n is the sze.le Standard deviation of these ages by usingthe computing formi..! -n-1h, ages by using the computing formula s the sample size. he standard deviation is Round to one decimal place as needed.) Compare your work in parts (b) and (c). answer to part (b) is (1)the answer to part (c). O the same as different from antitative data set under consideration has roughly a bell-shaped distribution. Apply the empirical ru
Answer #1

(a)

Range = Maximum - Minimum = 58 - 9 = 49

So,

Answer is:

49.

(b)

Sample mean (ar{x}) is given by:

ΣΕΙΖ, 160

x x - ar{x} (x -ar{x})2
21 -11 121
58 26 676
18 - 14 196
54 22 484
9 - 23 529
Total = 2006

So,

Standard Deviation is given by:

Σ51(ri-i)2 = /2006 = V501.5 = 22.4

So,

Answer is:

22.4

(c)

Sample mean (ar{x}) is given by:

ΣΕΙΖ, 160

x x2
21 441
58 3364
18 324
54 2916
9 81
Total = 160 7126

So,

Standard Deviation is given by:

71 n- rl

7126 1602 5-1 up /7126-5120-1200-= /501,5 = 22.4

So,

Answer is:

22.4

(d)

Answer to part (b) is same as the answer to part (c).